# Encoding and Controlling Global Semantics for Long-form Video Question Answering

### v2

@@ Abstract
Seeking answers effectively for long videos is essential to build video question answering (videoQA) systems. Previous methods adaptively select frames and regions from long videos to save computations. However, this fails to reason over the whole sequence of video, leading to sub-optimal performance. To address this problem, we introduce a state space layer (SSL) into multi-modal Transformer to efficiently integrate global semantics of the video, which mitigates the video information loss caused by frame and region selection modules. Our SSL includes a gating unit to enable controllability over the flow of global semantics into visual representations. To further enhance the controllability, we introduce a cross-modal compositional congruence (C3) objective to encourage global semantics aligned with the question. To rigorously evaluate long-form videoQA capacity, we construct two new benchmarks Ego-QA and MAD-QA featuring videos of considerably long length, i.e. 17.5 minutes and 1.9 hours, respectively. Extensive experiments demonstrate the superiority of our framework on these new as well as existing datasets.
@@ 2 Related Work
Cross-modal alignment. Alignment approaches, such as contrastive learning (Nguyen and Luu, 2021; Wu et al., 2023a; Nguyen et al., 2024e; Wu et al., 2024a), optimal transport (Nguyen and Luu, 2022; Wu et al., 2024d, c, 2023b), tree-based fusion (Nguyen et al., 2023c), and energy-based modeling (Nguyen et al., 2023d), have been investigated to compare different input samples. For multimodal learning, recent cross-modal alignment methods focus on establishing a shared latent space in which samples of different modalities can be compared readily (Nguyen et al., 2023b, 2024b, 2024d, 2024a, a, 2022b, 2022a). For example, with a contrastive learning formulation, CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) learn generalizable image-text representations from millions of image-text pairs. To exploit relation for cross-modal alignment, Ren et al. (2021) use hard bijective correspondences between words and objects via an argmax over the cross-modal attention matrix to optimize the cross-modal alignment. Pandey et al. (2022) generalize the hard alignment with their soft alignment approach, but they focus on objects within an image rather than a long video.
@@ 5.3 Baselines
We compare against the following state-of-the-arts: (i) HQGA (Xiao et al., 2022), which models video as a conditional graph hierarchy; (ii) CoVGT (Xiao et al., 2023), which exploits a graph transformer to encode video, along with contrastive learning to align its graph encoding with the textual encoding; (iii) ATP (Buch et al., 2022), which learns to select one salient video frame to perform videoQA; (iv) AIO (Wang et al., 2023), an end-to-end videoQA model that releases the need of unimodal encoders and leverages non-parametric token rolling operation; (v) STAGE (Lei et al., 2020), which jointly carries out videoQA with moment localization and object grounding; (vi) VQA-T (Yang et al., 2021), a cross-modal Transformer which is pre-trained with supervised contrastive learning on a large-scale videoQA dataset; (vii) RESERVE-B (Zellers et al., 2022), which learns to predict the masked textual and audio tokens given video frames; (viii) Flamingo-9B (Alayrac et al., 2022), which is able to process multi-modal prompt and cast videoQA as text prediction; (ix) EgoVLP (Lin et al., 2022a), an approach to pre-train on egocentric videos for egocentric videoQA; (x) VideoReCap (Islam et al., 2024), a recursive model that synergizes different video hierarchies to process hour-long videos; (xi) MIST (Gao et al., 2023), a multi-modal Transformer which decomposes video into frames, patches, and segments to efficiently process with self-attention mechanism.

### v3

@@ Abstract
Seeking answers effectively for long videos is essential to build video question answering (videoQA) systems. Previous methods adaptively select frames and regions from long videos to save computations. However, this fails to reason over the whole sequence of video, leading to sub-optimal performance. To address this problem, we introduce a state space layer (SSL) into multi-modal Transformer to efficiently integrate global semantics of the video, which mitigates the video information loss caused by frame and region selection modules. Our SSL includes a gating unit to enable controllability over the flow of global semantics into visual representations. To further enhance the controllability, we introduce a cross-modal compositional congruence (C3) objective to encourage global semantics aligned with the question. To rigorously evaluate long-form videoQA capacity, we construct two new benchmarks Ego-QA and MAD-QA featuring videos of considerably long length, i.e. 17.5 minutes and 1.9 hours, respectively. Extensive experiments demonstrate the superiority of our framework on these new as well as existing datasets. The code, model, and data have been made available at nguyentthong.github.io/Long_form_VideoQA.
@@ 2 Related Work
Cross-modal alignment. Alignment approaches, such as contrastive learning (Nguyen and Luu, 2021; Nguyen et al., 2024e; Wu et al., 2023a, 2024a), optimal transport (Nguyen and Luu, 2022; Wu et al., 2023b, 2024d, 2024c), tree-based fusion (Nguyen et al., 2023c), and energy-based modeling (Nguyen et al., 2023d), have been investigated to compare different input samples. For multimodal learning, recent cross-modal alignment methods focus on establishing a shared latent space in which samples of different modalities can be compared readily (Nguyen et al., 2023b, 2024b, 2024d, 2024a, a, 2022b, 2022a). For example, with a contrastive learning formulation, CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) learn generalizable image-text representations from millions of image-text pairs. To exploit relation for cross-modal alignment, Ren et al. (2021) use hard bijective correspondences between words and objects via an argmax over the cross-modal attention matrix to optimize the cross-modal alignment. Pandey et al. (2022) generalize the hard alignment with their soft alignment approach, but they focus on objects within an image rather than a long video.
@@ 5.3 Baselines
We compare our proposed framework against the following baseline:
  • HQGA (Xiao et al., 2022): models video as a conditional graph hierarchy which aggregates visual facts in a level-wise manner with the guidance of textual modality,
  • CoVGT (Xiao et al., 2023): a model which exploits a graph transformer to encode video by capturing visual objects and their relations for spatio-temporal reasoning, along with contrastive learning to align its graph transformer with the textual encoder,
  • ATP (Buch et al., 2022): a model which learns to select one salient video frame to perform the videoQA task,
  • AIO (Wang et al., 2023): an end-to-end videoQA model that releases the need of unimodal encoders and leverages non-parametric token rolling operation,
  • STAGE (Lei et al., 2020): a model that jointly carries out videoQA with moment localization and object grounding,
  • VQA-T (Yang et al., 2021): a cross-modal Transformer which is pre-trained with supervised contrastive learning on a large-scale videoQA dataset,
  • RESERVE-B (Zellers et al., 2022): jointly represents video frames, texts, and audio, and learns to predict the masked textual and audio tokens given video frames,
  • Flamingo-9B (Alayrac et al., 2022): a model which is able to process multi-modal prompt and cast videoQA as text prediction,
  • EgoVLP (Lin et al., 2022a): pre-trains multi-modal Transformer on egocentric videos for egocentric videoQA.
  • VideoReCap (Islam et al., 2024): a recursive model that synergizes different video hierarchies to process hour-long videos,
  • MIST (Gao et al., 2023): a multi-modal Transformer which decomposes video into frames, patches, and segments to efficiently process with its self-attention mechanism.